OpenStudy (anonymous):

1.) simplify the radical √96 2.) Express the radical in simplest form √121a^3b^2

OpenStudy (anonymous):

Answer choices for #1 : 16√2 4√6 8√12 Answer choices for #2 : 11ab√a 11ab 11ab√11ab

OpenStudy (ajprincess):

96=16*6 sqrt16=4 Therefore sqrt96=4sqrt6 sqrt121=11 sqrta^2=a sqrtb^2=b sqrt121a^3b^2=11absqrta

OpenStudy (ajprincess):

That is for 1st option b for 2nd option a